OpenStudy (anonymous):

integration of rational functions: Evaluate I=integral of (2x^2+3x+4)/(x^2+6x+10)dx

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ (2x^2+3x+4)/(x^2+6x+10) }dx\]

OpenStudy (amistre64):

most likely your gonna have to use partial fractions to split this up into do-able sums

OpenStudy (anonymous):

is there any special formula??

OpenStudy (amistre64):

like a magic trick? no but i do notice that the bottom poly appears to be prime; try to do division on it first to see how that plays out

OpenStudy (anonymous):

ax+b[d/dx(x)}+constant

OpenStudy (anonymous):

i'm doing correspondance cuorse sir, pls hlp

OpenStudy (amistre64):

i am helping, but i will not do it FOR you. what does the long division give us?

OpenStudy (amistre64):

if need be, try using the wolf to simplify it or dbl chk your results

OpenStudy (anonymous):

long divisin method?

OpenStudy (anonymous):

can we expand (x^2+6x+10)= (x+3)^2+1

OpenStudy (amistre64):

yes, but that will be more useful to us after we split this thing up into a more usable form.

OpenStudy (amistre64):

we can try to split it apart into individual sums as is and see how that plays out

OpenStudy (anonymous):

long divisn = 1+ (x^2-3x-6)/x^2+6x+10

OpenStudy (anonymous):

is ths correct?

OpenStudy (amistre64):

\[\frac{(2x^2+3x+4)}{(x^2+6x+10)}=\frac{2x^2}{(x+3)^2+1}+\frac{3x}{(x+3)^2+1}+\frac{4}{(x+3)^2+1}\]

OpenStudy (amistre64):

your long division seems off to me

OpenStudy (amistre64):

2 ---------------- x^2+6x+10 ( 2x^2+3x+4 (2x^2+12x+20) -------------- -9x-16 \[2-\frac{9x+16}{(x+3)^2+1}\]

OpenStudy (amistre64):

\[\int 2-\frac{9x}{(x+3)^2+1}-\frac{16}{(x+3)^2+1}\ dx\] the first term is simple enough, the middle term looks by partable, and the last is an arctan

OpenStudy (anonymous):

2x-9/2 tan^{-1} (x+3)-16 (tan^{-1}(x+3)+c

OpenStudy (anonymous):

is ths correct?

OpenStudy (amistre64):

first terms is good the middle and last term i think need more work

OpenStudy (amistre64):

last term is fine

OpenStudy (anonymous):

2x-9/2 x^{2} tan^{-1} (x+3)-16 tan^{-1}(x+3) +c

OpenStudy (amistre64):

-9/2 x^{2} tan^{-1} (x+3) this part is giving you troubles

OpenStudy (anonymous):

is ths correct

OpenStudy (anonymous):

thank you sooooo much sir

OpenStudy (amistre64):

\[\int \frac{9x}{(x+3)^2+1}dx =\frac{9}{2}ln((x+3)^2+1)-6tan^{-1}(x+3)\]

OpenStudy (anonymous):

but integral of 1/ (1+x^{2})= tan {-1} x

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=integrate+%282x%5E2%2B3x%2B4%29%2F%28x%5E2%2B6x%2B10%29 all in all, it should get to this

OpenStudy (anonymous):

thank you sir